Consider the following information for a hypothetical economy: C= 200+0.75(Y-T) I=G=250 T=200 where, Y=C + I + G All figures are in millions except the measures of responsiveness. A. What is the value of marginal propensity to save(mps)?(2pt) B. Find the equilibrium level of national income (2pts) C. Find the government expenditure multiplier? Using this multiplier, find the effect of 50 million increases in government purchase on equilibrium income? (2 pts)


1
Expert's answer
2021-09-23T17:37:22-0400

a)

MPS= 0.25

consumption function "= a+b(Y-T)" where a= autonomous Consumption and "b = MPC"

"MPS= 1-MPC = 1-0.75= 0.25"


b)

"Y=C + I + G"


"Y = 200+0.75(Y-300) +250+250"


"Y= 200+0.75Y - 225+500"


"Y-0.75Y = 475"


"0.25Y = 475"


"Y= 1900\space million"


c)

Government multiplier "= \frac{1}{MPS} = \frac{1}{0.25}= 4"


When government spending increase by $50 million , Equilibrium income increase "= 4\times50 = 200" million